# Robust Learning of Multi-index Models via Iterative Subspace Approximation

### v1

@@ 3.1 Finding a Relevant Direction
In this section, we prove Proposition 3.7, which states that if the class satisfies Definition 3.2 and the current approximation of the ground-truth subspace provides a suboptimal piecewise approximation, then Algorithm 4 outputs a small set of vectors, at least one of which has a non-trivial projection onto the ground-truth subspace.
@@
Let $d,k\in\mathbb{Z}_{+}$, $\epsilon,\delta\in(0,1)$ and $\epsilon^{\prime}{\leq}c\epsilon^{3}/(\lvert\mathcal{Y}\rvert\Gamma^{2}k^{2}))$, for $c$ a sufficiently small universal constant. Let $D$ be a distribution supported on $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, let $V\subseteq\mathbb{R}^{d}$ be a $k$-dimensional subspace, and let $\mathcal{F}$ be a class of $({m,\zeta+O(\epsilon)},\alpha,K,\tau,\sigma,\Gamma)$-well-behaved Multi-Index Models according to Definition 3.2. Let $f\in\mathcal{F}$ with $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]={\mathrm{OPT}}$ and let $W$ be a subspace such that $f(\mathbf{x})=f(\mathbf{x}^{W})$ for all $\mathbf{x}\in\mathbb{R}^{d}$, with $K=\dim(W)$. Additionally, let $h_{\mathcal{S}}$ be a piecewise constant approximation of $D$, with respect to an $\epsilon^{\prime}$-approximating partition $\mathcal{S}$, according $V$, as defined in Definitions 2.2 and 2.3.
There exists $N=(dm)^{O(m)}{(k/\epsilon^{\prime})^{O(k)}}\log(1/\delta){(K\Gamma\lvert\mathcal{Y}\rvert/(\sigma\epsilon\alpha))^{O(1)}}$ such that, If $\zeta\geq\mathrm{OPT}$ and $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]>{\tau+\mathrm{OPT}+3\epsilon}$, then Algorithm 4 given $N$ i.i.d. samples from $D$, runs in $\mathrm{poly}(N,d,\lvert\mathcal{Y}\rvert)$ time, and with probability at least $1-\delta$, returns a list of unit vectors $\mathcal{E}$ of size $|\mathcal{E}|=\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))$, such that: For some $\mathbf{v}\in\mathcal{E}$ and unit vector $\mathbf{w}\in W$ it holds that
@@
Before proving Proposition 3.7, we need to show that if we have learned a direction of our target model to accuracy $\epsilon$, then we only observe moments that depend on the remaining directions. The intuition behind the proof is that re-randomizing over directions that have negligible projection to $W$ barely affects the function’s output. Hence, the function is close to a mixing $y^{\prime}$ of the noisy version $y$ that has these components re-randomized. As a result, if the function is far from a projection onto $V$, we have large moments for level-sets of $y^{\prime}$. Consequently, the re-randomized directions are averaged out, and we are left with distinguishing moments on the remaining directions. The proof is similar in structure to that of Lemma 3.5, the main difference being in what is being re-randomized.
@@
Let $D$ be a distribution over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, let $\mathcal{F}$ be a class of $(m,\epsilon+O(\epsilon^{\prime}),\delta,K,\tau,\sigma,\Gamma)$-well-behaved Multi-Index Models according to Definition 3.2. Let $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$, with $k\geq 1$, and let $\mathcal{S}$ be an $\eta$-approximating partition of $V$ according to Definition 2.2, with $\eta=\epsilon^{\prime 3}/(C\lvert\mathcal{Y}\rvert\Gamma^{2}k^{2})$, for $C>0$ a sufficiently large universal constant. Let $f\in\mathcal{F}$ be a function that depends only on a $K$-dimensional subspace $W$, and let $y$ be random variable over $\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]\leq\epsilon$.
Suppose that there exists no function $g:V\to\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\leq\tau$. Then, there exists $\mathcal{T}\subseteq\mathcal{S}$ with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}[S]\geq\delta$ such that for each $S\in\mathcal{T}$ there exists a zero-mean, variance-one polynomial $p:U\to\mathbb{R}$ of degree at most $m$ such that $\operatorname*{\mathbf{E}}_{\mathbf{x},y}[p(\mathbf{x}^{U})\mathds{1}(y=z)\mid\mathbf{x}\in S]>\sigma$, where $U{=}(V{+}W)\cap V^{\perp}$. Moreover, for each such polynomial $p$, we have that $\nabla p(\mathbf{x}^{U})\cdot\mathbf{v}=0$ for any $\mathbf{v}\in E$ and $\mathbf{x}\in\mathbb{R}^{d}$, where $E\subseteq{V^{\perp}}$ is a subspace satisfying $\|\mathbf{v}^{W}\|\leq\epsilon^{\prime}/\Gamma(f)$ for every unit vector $\mathbf{v}\in E$.
@@
Let $\mathbf{x},\mathbf{x}^{\prime}$ be random variables defined as $\mathbf{x}=\mathbf{x}^{\perp E}+\mathbf{x}^{E}$ and $\mathbf{x}^{\prime}=\mathbf{x}^{\perp E}+{\mathbf{z}}^{E}$, where $\mathbf{z}$ is an independent standard Gaussian vector. Denote by $\mathcal{S}$ an $\eta$-approximating partition with respect to $V$. Notice that for any unit direction $\mathbf{w}\in W$, both $\mathbf{w}\cdot\mathbf{x}$ and $\mathbf{w}\cdot\mathbf{x}^{\prime}$ are distributed according to a standard Gaussian. Furthermore, we have that by independence
@@
Given that $E$ is the subspace of small projections, i.e., $\left\|\mathbf{w}^{E}\right\|\leq\epsilon^{\prime}/\Gamma(f)$; it follows that $\operatorname*{\mathbf{E}}[(\mathbf{w}\cdot\mathbf{x}^{\perp E})^{2}]\geq 1-(\epsilon^{\prime}/\Gamma(f))^{2}$. As a result, $\mathbf{x}$ and $\mathbf{x}^{\prime}$ are at least $1-(\epsilon^{\prime}/\Gamma(f))^{2}$ correlated in $W$. Thus by the Gaussian noise sensitivity bound A.5, we have that
$\operatorname*{\mathbf{Pr}}_{\mathbf{x},\mathbf{x}^{\prime}}[f(\mathbf{x})\neq f(\mathbf{x}^{\prime})]\leq O(\epsilon^{\prime})\;.$
Consider the random variable $y^{\prime}$ supported on $\mathcal{Y}$ that is distributed like $y$ for $\mathbf{x}^{\prime}$, i.e., $D(y^{\prime}=i\mid\mathbf{x}^{\prime}=\mathbf{z})=D(y=i\mid\mathbf{x}=\mathbf{z})$, for all $\mathbf{z}\in\mathbb{R}^{d},i\in\mathcal{Y}$. We have that
$\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y^{\prime}]\leq\operatorname*{\mathbf{Pr}}[f(\mathbf{x}^{\prime})\neq y^{\prime}]+O(\epsilon^{\prime})\leq\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y]+O(\epsilon^{\prime})\leq\epsilon+O(\epsilon^{\prime})\;.$
As a result, by Lemma 3.5, we have that there exists $\mathcal{T}\subseteq\mathcal{S}$ with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}[S]\geq\delta$ such that for each $S\in\mathcal{T}$ there exists a zero-mean, variance-one polynomial $p:U\to\mathbb{R}$ of degree at most $m$ along with a label $z\in\mathcal{Y}$ such that $\operatorname*{\mathbf{E}}_{\mathbf{x},y^{\prime}}[p(\mathbf{x}^{U})\mathds{1}(y^{\prime}=z)\mid\mathbf{x}\in S]>\sigma$, where $U=(V+W)\cap V^{\perp}$. However we have that for all $S\in\mathcal{T}$
$\operatorname*{\mathbf{E}}_{\mathbf{x},y^{\prime}}[p(\mathbf{x}^{U})\mathds{1}(y^{\prime}=z)\mid\mathbf{x}\in S]=\operatorname*{\mathbf{E}}[\operatorname*{\mathbf{E}}_{\mathbf{x}^{E}}[p(\mathbf{x}^{U})\mid\mathbf{x}^{\perp E}]\operatorname*{\mathbf{E}}_{\mathbf{x}^{E}}[\mathds{1}(y=i)\mid\mathbf{x}^{\perp E}]\mid\mathbf{x}\in S]]\;.$
Notice that $p^{\prime}(\mathbf{x}^{U})\coloneqq\operatorname*{\mathbf{E}}_{\mathbf{x}^{E}}[p(\mathbf{x}^{U})\mid\mathbf{x}^{\perp E}]$ is a mean-zero polynomial of degree at most $m$ with variance at most one by Jensen’s inequality. Furthermore, as $p^{\prime}$ is independent of $\mathbf{x}^{E}$ we have that $\nabla p^{\prime}(\mathbf{x}^{U})\cdot\mathbf{v}=0$ for any $\mathbf{v}\in E$ and $\mathbf{x}\in\mathbb{R}^{d}$. ∎
@@
Denote by $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}$ an orthonormal basis of the subspace $W$. Also, let $\widehat{\mathbf{U}}$ be the matrix computed at Line 6 of Algorithm 4 and let $\eta^{2}$ be the $l_{2}^{2}$ error chosen in the polynomial-regression step, i.e., Line 4 of Algorithm 4.
First note that as $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]>{\tau+\mathrm{OPT}+3\epsilon}$ we have that form Lemma 3.4 for every function $g:V\to\mathcal{Y}$ it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\geq\tau$. Thus from Lemma 3.8 we have that there exists $\mathcal{T}\subseteq\mathcal{S}$ with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}[S]\geq\alpha$, such that for each $S\in\mathcal{T}$ there exists a $z\in\mathcal{Y}$ and a zero-mean variance-one polynomial $q_{S}:U\to\mathbb{R}$ of degree at most $m$ such that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[q_{S}(\mathbf{x}^{U})\mathds{1}(y=z)\mid\mathbf{x}\in S]>\sigma$, where $U=(V+W)\cap(V^{\perp})$. Furthermore, define $E_{S}\coloneqq\mathrm{span}(\{\mathbf{v}\in U_{S}:\left\|\mathbf{v}^{W}\right\|\leq\epsilon/(\sqrt{\lvert U_{S}\rvert}\Gamma)\})$. Then, for every unit vector $\mathbf{v}\in E_{S}$, we have $\nabla q_{S}(\mathbf{x}^{U})\cdot\mathbf{v}=0$ since $\left\|\mathbf{v}^{W}\right\|\leq\epsilon/\Gamma$ holds for every unit vector $\mathbf{v}\in E_{S}$.
In the following claim, we leverage the existence of non-trivial moments over $U$ for regions $S\in\mathcal{T}$, to show that if the sample size is sufficiently large then for each region $S\in\mathcal{T}$ the associated set, $U_{S}$ (see Line 5 of Algorithm 4) contains a vector that correlates with $W$.
@@
Fix, $S\in\mathcal{T}$, if the number of samples that fall in $S$ is $N_{S}\geq(dm)^{Cm}\log(\lvert\mathcal{Y}\rvert/\delta)/{\eta}^{C}$, for a sufficiently large universal constant $C>0$, then with probability at least $1-\delta$ there exists a unit vector $\mathbf{v}\in U_{S}$ such that
$(\mathbf{w}^{(i)})^{\perp V}\cdot\mathbf{v}\geq{\mathrm{poly}(\epsilon\sigma/(mK\lvert\mathcal{Y}\rvert\Gamma)}\;,$
for some $i\in[K]$. Moreover, it holds that $\lvert U_{S}\rvert=\mathrm{poly}(mK\lvert\mathcal{Y}\rvert/\sigma)$.
@@
Denote by $p_{S,i}$ for $S\in\mathcal{S},i\in\mathcal{Y}$ the polynomials of degree at most $m$ computed at Line 4 of Algorithm 4. First, notice that since whether or not $\mathbf{x}\in S$ depends only in $\mathbf{x}^{V}$, we have that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}^{\perp V}))^{2}\mid\mathbf{x}\in S]\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}^{\perp V}))^{2}\mid\mathbf{x}\in S]+\eta^{2}$ is equivalent to $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}))^{2}]\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}))^{2}]+\eta^{2}$, where $D_{V^{\perp}}$ is defined to be the marginal obtained by averaging $D$ over $V$ conditioned on $S$, i.e., $D_{V^{\perp}}(\mathbf{x}^{\perp V},y)=\operatorname*{\mathbf{E}}_{\mathbf{x}^{V}}[D(\mathbf{x},y)\mid\mathbf{x}\in S]$. Moreover, by the properties of the Gaussian distribution, we have that the $\mathbf{x}$-marginal of $D_{V^{\perp}}$ is a standard Gaussian. Therefore, if $N_{S}\geq(dm)^{Cm}\log(\lvert\mathcal{Y}\rvert/\delta)/{\eta^{C}}$ for a sufficiently large universal constant $C>0$, then by the union bound and B.2, we have that with probability at least $1-\delta$ it holds that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}))^{2}]\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}))^{2}]+\eta^{2}$. Moreover, by the orthogonality of Hermite polynomials we have that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}))^{2}]=\sum_{\beta\subseteq\mathbb{N}^{d}}(\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[H_{\beta}(\mathbf{x})\mathds{1}(y=i)]-\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[H_{\beta}(\mathbf{x})p_{S,i}(\mathbf{x})])^{2}$, it holds that $\sum_{t=1}^{m}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)^{[t]}-p_{S,i}^{[t]}(\mathbf{x}))^{2}]\leq\eta^{2}$.

### v2

@@ 3.1 Finding a Relevant Direction
In this section, we prove Proposition 3.7, which states that if the target function class satisfies Definition 3.2 and the current approximation of the ground-truth subspace provides a suboptimal piecewise approximation, then Algorithm 4 outputs a small set of vectors, at least one of which has a non-trivial projection onto the ground-truth subspace.
@@
Let $d,k\in\mathbb{Z}_{+}$, $\epsilon,\delta\in(0,1)$ and $\epsilon^{\prime}\leq c\epsilon^{3}/(\lvert\mathcal{Y}\rvert\Gamma^{2}k^{2}))$, for $c$ a sufficiently small universal constant. Let $D$ be a distribution supported on $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, let $V\subseteq\mathbb{R}^{d}$ be a $k$-dimensional subspace, and let $\mathcal{F}$ be a class of $(m,\zeta+O(\epsilon),\alpha,K,\tau,\sigma,\Gamma)$-well-behaved Multi-Index Models according to Definition 3.2. Let $f\in\mathcal{F}$ with $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]=\mathrm{OPT}$ and let $W$ be a subspace such that $f(\mathbf{x})=f(\mathbf{x}^{W})$ for all $\mathbf{x}\in\mathbb{R}^{d}$, with $K=\dim(W)$. Additionally, let $h_{\mathcal{S}}$ be a piecewise constant approximation of $D$, with respect to an $\epsilon^{\prime}$-approximating partition $\mathcal{S}$, according to $V$, as defined in Definitions 2.2 and 2.3.
There exists $N=(dm)^{O(m)}(k/\epsilon^{\prime})^{O(k)}\log(1/\delta)(K\Gamma\lvert\mathcal{Y}\rvert/(\sigma\epsilon\alpha))^{O(1)}$ such that: If $\zeta\geq\mathrm{OPT}$ and $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]>\tau+\mathrm{OPT}+3\epsilon$, then Algorithm 4 given $N$ i.i.d. samples from $D$, runs in $\mathrm{poly}(N)$ time, and with probability at least $1-\delta$, returns a list of unit vectors $\mathcal{E}$ of size $|\mathcal{E}|=\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))$ such that: For some $\mathbf{v}\in\mathcal{E}$ and unit vector $\mathbf{w}\in W$ it holds that
@@
Before proving Proposition 3.7, we show that once a direction of our target model is learned to sufficiently small accuracy, the observed moments depend only on the remaining directions.
The intuition is as follows. Re-randomizing over directions that have negligible projection onto $W$ barely affects the output of the function. Thus, the function is close to a version $y^{\prime}$ of the noisy label $y$ in which these components have been averaged. Consequently, if the function is far from being a function of the current subspace $V$ (i.e., it still depends significantly on directions outside $V$), then the level sets of $y^{\prime}$ will exhibit large distinguishing moments. In this way, the influence of the re-randomized (negligible) directions is eliminated, leaving only the significant, distinguishing moments on the remaining directions.
The proof follows a structure similar to that of Lemma 3.5, with the main difference being the set of directions over which re-randomization occurs.
@@
Let $D$ be a distribution over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$. Let $\mathcal{F}$ be a class of $(m,\zeta+O(\epsilon),\alpha,K,\tau,\sigma,\Gamma)$-well-behaved Multi-Index Models as defined in Definition 3.2. Let $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$, with $k\geq 1$, and let $\mathcal{S}$ be an $\eta$-approximating partition of $V$ (see Definition 2.2), where $\eta=\epsilon^{3}/(C\lvert\mathcal{Y}\rvert\Gamma^{2}k^{2})$ and $C>0$ is a sufficiently large universal constant. Let $f\in\mathcal{F}$ be a function that depends only on a $K$-dimensional subspace $W$, and let $y$ be random variable over $\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]\leq\zeta$.
Suppose that no function $g:V\to\mathcal{Y}$ exists with $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\leq\tau$. Then, there exists $\mathcal{T}\subseteq\mathcal{S}$ with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}[S]\geq\alpha$ such that for each $S\in\mathcal{T}$ the following hold:
  1. 1.
    There exists a label $z\in\mathcal{Y}$ and a zero-mean, variance-one polynomial $p:U\to\mathbb{R}$ of degree at most $m$ such that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[p(\mathbf{x}^{U})\mathds{1}(y=z)\mid\mathbf{x}\in S]>\sigma$, where $U{=}(V{+}W)\cap V^{\perp}$.
  2. 2.
    For each such polynomial $p$ satisfying (1), it holds that $\nabla p(\mathbf{x}^{U})\cdot\mathbf{v}=0$ for every $\mathbf{v}\in E$ and all $\mathbf{x}\in\mathbb{R}^{d}$, where $E\subseteq{V^{\perp}}$ is a subspace satisfying $\|\mathbf{v}^{W}\|\leq\epsilon/\Gamma(f)$ for every unit vector $\mathbf{v}\in E$.
@@
Let $\mathbf{x},\mathbf{x}^{\prime}$ be random variables defined as $\mathbf{x}=\mathbf{x}^{\perp E}+\mathbf{x}^{E}$ and $\mathbf{x}^{\prime}=\mathbf{x}^{\perp E}+{\mathbf{z}}^{E}$, where $\mathbf{z}$ is an independent standard Gaussian vector. Denote by $\mathcal{S}$ an $\eta$-approximating partition with respect to $V$. Notice that for any unit direction $\mathbf{w}\in W$, both $\mathbf{w}\cdot\mathbf{x}$ and $\mathbf{w}\cdot\mathbf{x}^{\prime}$ are distributed according to a standard Gaussian. Furthermore, by independence we have that
@@
Given that $E$ is the subspace of small projections, (i.e., for any unit vector $\mathbf{w}\in W$, we have $\|\mathbf{w}^{E}\|\leq\epsilon/\Gamma(f)$), it follows that $\operatorname*{\mathbf{E}}[(\mathbf{w}\cdot\mathbf{x}^{\perp E})^{2}]\geq 1-(\epsilon/\Gamma(f))^{2}$ for any unit vector $\mathbf{w}\in W$. This means that $\mathbf{x}$ and $\mathbf{x}^{\prime}$ are at least $1-(\epsilon/\Gamma(f))^{2}$ correlated in $W$. By the Gaussian noise sensitivity bound (see A.5), we have that
$\operatorname*{\mathbf{Pr}}_{\mathbf{x},\mathbf{x}^{\prime}}[f(\mathbf{x})\neq f(\mathbf{x}^{\prime})]\leq O(\epsilon)\;.$
Next, define a new random variable $y^{\prime}$ over $\mathcal{Y}$ which is distributed like $y$ but with respect to $\mathbf{x}^{\prime}$; that is, for every $\mathbf{z}\in\mathbb{R}^{d}$ and $i\in\mathcal{Y}$, $D(y^{\prime}=i\mid\mathbf{x}^{\prime}=\mathbf{z})=D(y=i\mid\mathbf{x}=\mathbf{z})$. We have that
$\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y^{\prime}]\leq\operatorname*{\mathbf{Pr}}[f(\mathbf{x}^{\prime})\neq y^{\prime}]+O(\epsilon)\leq\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y]+O(\epsilon)\leq\zeta+O(\epsilon)\;.$
As a result, by Lemma 3.5, we have that there exists $\mathcal{T}\subseteq\mathcal{S}$ with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}[S]\geq\alpha$ such that for each $S\in\mathcal{T}$ there exists a zero-mean, variance-one polynomial $p:U\to\mathbb{R}$ of degree at most $m$ and a label $z\in\mathcal{Y}$ such that $\operatorname*{\mathbf{E}}_{\mathbf{x},y^{\prime}}[p(\mathbf{x}^{U})\mathds{1}(y^{\prime}=z)\mid\mathbf{x}\in S]>\sigma$, where $U=(V+W)\cap V^{\perp}$. However we have that for all $S\in\mathcal{T}$
$\operatorname*{\mathbf{E}}_{\mathbf{x},y^{\prime}}[p(\mathbf{x}^{U})\mathds{1}(y^{\prime}=z)\mid\mathbf{x}\in S]=\operatorname*{\mathbf{E}}\left[\operatorname*{\mathbf{E}}_{\mathbf{x}^{E}}[p(\mathbf{x}^{U})\mid\mathbf{x}^{\perp E}]\operatorname*{\mathbf{E}}_{\mathbf{x}^{E}}[\mathds{1}(y=i)\mid\mathbf{x}^{\perp E}]\mid\mathbf{x}\in S\right]\;,$
where we used that $y^{\prime}$ is independent of $\mathbf{x}^{E}$. Notice that $p^{\prime}(\mathbf{x}^{U})\coloneqq\operatorname*{\mathbf{E}}_{\mathbf{x}^{E}}[p(\mathbf{x}^{U})\mid\mathbf{x}^{\perp E}]$ is a zero-mean polynomial of degree at most $m$ with variance at most one by Jensen’s inequality. Furthermore, as $p^{\prime}$ is independent of $\mathbf{x}^{E}$ we have that $\nabla p^{\prime}(\mathbf{x}^{U})\cdot\mathbf{v}=0$ for any $\mathbf{v}\in E$ and $\mathbf{x}\in\mathbb{R}^{d}$. ∎
@@
Denote by $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}$ an orthonormal basis for the subspace $W$. Let $\widehat{\mathbf{U}}$ be the matrix computed in Line 6 of Algorithm 4 and let $\eta^{2}$ be the $L_{2}^{2}$ error chosen in the polynomial-regression step (see Line 4 of Algorithm 4).
First, observe that since $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]>\tau+\mathrm{OPT}+3\epsilon$, Lemma 3.4 for every implies that for every function $g:V\to\mathcal{Y}$ it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\geq\tau$. Then, by Lemma 3.8 there exists a $\mathcal{T}\subseteq\mathcal{S}$ with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}[S]\geq\alpha$, such that for each $S\in\mathcal{T}$ there exists a label $z\in\mathcal{Y}$ and a zero-mean variance-one polynomial $q_{S}:U\to\mathbb{R}$ of degree at most $m$ satisfying $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[q_{S}(\mathbf{x}^{U})\mathds{1}(y=z)\mid\mathbf{x}\in S]>\sigma$, where $U=(V+W)\cap(V^{\perp})$. Furthermore, for each $S\in\mathcal{S}$, define $E_{S}\coloneqq\mathrm{span}(\{\mathbf{v}\in U_{S}:\left\|\mathbf{v}^{W}\right\|\leq\epsilon/(\sqrt{\lvert U_{S}\rvert}\Gamma)\})$. Then, for every unit vector $\mathbf{v}\in E_{S}$, we have $\nabla q_{S}(\mathbf{x}^{U})\cdot\mathbf{v}=0$ since $\left\|\mathbf{v}^{W}\right\|\leq\epsilon/\Gamma$ holds for every unit vector $\mathbf{v}\in E_{S}$.
In the following claim, we leverage the existence of non-trivial moments over $U$ for regions $S\in\mathcal{T}$. We show that if the sample size is sufficiently large then for each region $S\in\mathcal{T}$ the associated set $U_{S}$ (see Line 5 of Algorithm 4) contains a vector that correlates with $W$.
@@
Let $C>0$ be a sufficiently large universal constant. Fix $S\in\mathcal{T}$. If the number of samples that fall in $S$ is $N_{S}\geq(dm)^{Cm}\log(\lvert\mathcal{Y}\rvert/\delta)/\eta^{C}$, then with probability at least $1-\delta$ there exists a unit vector $\mathbf{v}\in U_{S}$ and $i\in[K]$ such that
$(\mathbf{w}^{(i)})^{\perp V}\cdot\mathbf{v}\geq\mathrm{poly}(\epsilon\sigma/(mK\lvert\mathcal{Y}\rvert\Gamma)\;.$
Moreover, it holds that $\lvert U_{S}\rvert=\mathrm{poly}(mK\lvert\mathcal{Y}\rvert/\sigma)$.
@@
Denote by $p_{S,i}$ (for $S\in\mathcal{S}$ and $i\in\mathcal{Y}$) the degree at most $m$ polynomial computed at Line 4 of Algorithm 4. Notice that membership of a point $\mathbf{x}$ in a cube $S$ depends solely on its projection onto $V$, i.e., on $\mathbf{x}^{V}$. Therefore, the error guarantee obtained in the regression step $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}^{\perp V}))^{2}\mid\mathbf{x}\in S]\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}^{\perp V}))^{2}\mid\mathbf{x}\in S]+\eta^{2}$ is equivalent to $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}))^{2}]\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}))^{2}]+\eta^{2}$, where $D_{V^{\perp}}^{S}$ is the marginal obtained by averaging $D$ over $V$ conditioned on $S$, i.e., $D^{S}_{V^{\perp}}(\mathbf{x}^{\perp V},y)=\operatorname*{\mathbf{E}}_{\mathbf{x}^{V}}[D(\mathbf{x},y)\mid\mathbf{x}\in S]$.
@@
Moreover, by the properties of the Gaussian distribution, we have that the $\mathbf{x}$-marginal of $D_{V^{\perp}}$ is a standard Gaussian. Therefore, if $N_{S}\geq(dm)^{Cm}\log(\lvert\mathcal{Y}\rvert/\delta)/\eta^{C}$ for a sufficiently large universal constant $C>0$, then by applying the union bound and B.2, we have that with probability at least $1-\delta$:
$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}))^{2}]\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}))^{2}]+\eta^{2}\;.$
Furthermore, by the orthogonality of Hermite polynomials, we have that
$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}))^{2}]=\sum_{\beta\subseteq\mathbb{N}^{d}}(\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[H_{\beta}(\mathbf{x})\mathds{1}(y=i)]-\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[H_{\beta}(\mathbf{x})p_{S,i}(\mathbf{x})])^{2}\;.$
In particular, if we decompose the error into its Hermite polynomial components of degree $t$, we have that $\sum_{t=1}^{m}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[(\mathds{1}(y=i)^{[t]}-p_{S,i}^{[t]}(\mathbf{x}))^{2}]\leq\eta^{2}$.